A student scored 70 points on a test that had 4 seven point fill in the blank questions and 24 three point multiple choice questions. If the student answered 1 fill in the blank questions correctly, how many multiple choice questions did the student answer correctly?
Answer by JoelSchwartz(130)

x=number of multiple choice questions the student got wrong.
7(4-3)+3(24-x)=70
7+3(24-x)=70
7+72-3x=70
79-3x=70
x=3